OpenStudy (pandagurl):

Solve. 4x – 6 = −3x + 8 x =

OpenStudy (pandagurl):

MEDAL

OpenStudy (yuliacosta111):

x=14

OpenStudy (pandagurl):

..... can you explain?

OpenStudy (yuliacosta111):

sure

OpenStudy (pandagurl):

its not 14,,, I tested it

OpenStudy (yuliacosta111):

really brb

OpenStudy (anonymous):

no X=2

OpenStudy (pandagurl):

4x14 - 6 = 50 -3x14 + 8 = -34

OpenStudy (anonymous):

you add 3x to 4x and add 6 to 8 and end up with 7x=14 and reduce

OpenStudy (pandagurl):

thank you! :D (tested it with my calculator and your correct) I just did not know where to start...

OpenStudy (anonymous):

your welcome and i did it in my head

OpenStudy (pandagurl):

whoa

OpenStudy (anonymous):

its actually not that hard